With a certain tail wind a jet aircraft arrives at its destination, 1,890 miles away, in 3 hours. Flying against the same wind, the plane makes the return trip in 3 3/8 hours. Find the wind speed and the plane's airspeed.
Answer by TimothyLamb(4379)

---
s = d / t
---
outbound:
s + w = 1890/3
s = 1890/3 - w
---
return:
s - w = 1890/(3+3/8)
s = 1890/(24/8+3/8) + w
s = 1890/(27/8) + w
---
s + w = 1890/3
s - w = 1890/(3+3/8)
---
put the system of linear equations into standard form
---
x = airplane airspeed
y = wind speed
---
x + y = 630
x - y = 560
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = airplane airspeed = 595 mph
y = wind speed = 35 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php